5/6 x 1/5 + 4 = 7/10 x
solve the equation and check for the solution

Answer by rfer(16322)

25/30+6/30+120/30=21/30x
(151/30)/(21/30)=x
151/30*30/21=x
x=4530/630
x=7 4/21